# A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability

### v2

@@ Abstract.
Crypto-assets and central bank digital currencies (CBDCs) are reshaping how value is exchanged in distributed computing environments. These systems combine cryptographic primitives, protocol design, and system architectures to provide transparency and efficiency while raising critical challenges around privacy and regulatory compliance. This survey offers a comprehensive overview of privacy-preserving digital payment systems, covering both decentralized ledgers and CBDCs. We present a taxonomy of privacy goals—including anonymity, confidentiality, unlinkability, and auditability—and map them to underlying cryptographic primitives, protocols, and system architectures. Our work adopts a design-oriented perspective, linking high-level privacy objectives to concrete implementations. We also trace the evolution of privacy-preserving digital payment systems through three generations, highlighting shifts from basic anonymity guarantees toward more nuanced privacy-accountability trade-offs. Finally, we identify open challenges, motivating further research into architectures and solutions that balance strong privacy with real-world auditability needs.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: DLTisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The rapid proliferation of digital payment systems has sparked renewed interest in designing distributed architectures that balance user privacy with institutional accountability. Spanning decentralized crypto-assets [^1] The term cryptocurrency is widely used to describe decentralized systems such as Bitcoin, Ethereum, or Monero, that enable peer-to-peer asset transfer without a central authority. However, this usage differs from the terminology adopted by central banks and regulators (ECB, 2018), which reserve the term “currency” for instruments with legal tender status. To reflect this distinction and bridge the terminology gap between these communities, we adopt the more nuanced terms crypto-assets or digital payment systems throughout the paper. and centrally managed infrastructures such as central bank digital currencies (CBDCs), these systems leverage cryptographic techniques and distributed consensus to enable secure, efficient, and programmable value transfer. However, the same transparency and auditability that make these systems attractive also raise fundamental questions about user privacy, institutional control, and regulatory compliance (e.g., (Auer et al., 2025; Ghesmati et al., 2022; Herskind et al., 2020)). For example, public blockchains typically offer pseudo-anonymity. This enables transactions tracing, address clustering, and ultimately users de-anonymization, thus exposing financial behavior (e.g., (Meiklejohn et al., 2016; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). As a result, privacy-preserving digital payments have emerged as a critical research field, attracting attention from cryptographers, engineers, economists, and policymakers.
@@
Researchers have developed a variety of privacy-enhancing technologies (PETs), including mixing protocols (e.g., (Maxwell, 2013; Ziegeldorf et al., 2018; Ruffing and Moreno-Sanchez, 2017)) and confidential transactions (e.g., (Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022)), and privacy-native crypto solutions (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Tomescu et al., 2022; Noether et al., 2016)). They implement various notions of privacy, such as anonymity, confidentiality, and unlinkability. However, privacy alone is not sufficient: compliance with regulatory frameworks—such as Anti-Money Laundering (AML) and Combating the Financing of Terrorism (CFT)—requires mechanisms for auditability and selective disclosure. Therefore, recent work has explored cryptographic protocols such as zero-knowledge proofs (ZKPs), threshold encryption, and homomorphic commitments to reconcile these conflicting goals.
@@
This survey proposes a unifying, design-oriented perspective on privacy-preserving digital payment systems, covering both decentralized ledgers and centrally managed systems with privacy guarantees—including CBDCs. Unlike previous surveys that focus on broad overviews (e.g., (Genkin et al., 2018; Auer et al., 2025; Herskind et al., 2020)), specific technologies (e.g., (Almashaqbeh and Solomon, 2022; Ghesmati et al., 2022; Sun et al., 2021)), or do not consider auditability (e.g., (Peng et al., 2021; Herskind et al., 2020; Feng et al., 2019)), our work aims to bridge these dimensions. Our contributions are threefold. First, we introduce a taxonomy that classifies privacy-preserving digital payments systems along multiple dimensions, including system architecture, trust model, value representation, privacy guarantees, and auditability mechanisms. Second, we systematize how privacy properties are formalized and realized, mapping high-level goals to technical mechanisms. We trace the development of privacy-preserving digital payment systems, highlighting the role of cryptographic techniques across three generations. Third, we identify open research challenges stemming from the design of privacy-preserving digital payment systems, which lie at the intersection of cryptography, distributed systems, and privacy engineering. Supplementary material further contextualizes key publication venues and technical details of specific privacy-enhancing mechanisms.
@@ 2. Related Surveys
Bonneau et al. (Bonneau et al., 2015) reviewed extensions of Bitcoin and introduced altcoins (i.e., alternatives to Bitcoin), including early privacy-native protocols. Genkin et al. (Genkin et al., 2018) offered a broad and accessible overview of techniques for anonymity in decentralized crypto solutions, while emphasizing the absence of a unified definition of privacy. More recent surveys attempted to formalize anonymity notions. For instance, Amarasinghe et al. (Amarasinghe et al., 2019) developed a taxonomy of anonymity features, and evaluated Bitcoin-derived solutions against it. Our work draws from and extends this analysis by further considering the trade-offs between anonymity, confidentiality, and auditability across a broader range of systems, including CBDCs. Feng et al. (Feng et al., 2019) offered a layered classification of privacy-preserving mechanisms from network-level obfuscation to cryptographic primitives. While technically detailed, their coverage of payment protocols and privacy-accountability trade-offs is limited. Similarly, a systematic literature review by Herskind et al. (Herskind et al., 2020) catalogs techniques such as stealth addresses, confidential transactions, and network anonymity. However, it does not analyze core system features such as value representation models or system architectures. Peng et al. (Peng et al., 2021) identify PETs and illustrate how various protocols integrate them, but do not investigate protocol evolution and design trade-offs.
@@
A distinct body of work addresses privacy in CBDCs, driven by the need to balance individual privacy with institutional oversight. Auer et al. (Auer et al., 2022, 2025) distinguish between soft privacy, based on access controls and trust in intermediaries, and hard privacy, based on cryptographic techniques. While they offer a high-level classification and conceptual framework, their discussion does not consider low-level protocol mechanisms. Darbha and Arora (Bank of Canada) [^2] https://www.bankofcanada.ca/2020/06/staff-analytical-note-2020-9/ and Pocher and Veneris (Pocher and Veneris, 2022) explore PETs from a compliance and governance perspective. These works offer blueprints but lack detailed cryptographic analysis. Conversely, our work presents a comparative review of technical design choices including primitives, protocols, and architectures.
@@
We propose a unified taxonomy of privacy-preserving payment systems that integrates both decentralized ledger-based solutions and CBDCs. Specifically, we present a systematic analysis of existing solutions to identify key value representation models, privacy features, and auditability guarantees. We focus on how such features are realized through cryptographic protocols and specific design choices, resulting in a classification of contributions that connects design goals to technical implementations.
@@ 3.1. Commitments and Homomorphic Commitments
A commitment scheme is a cryptographic primitive that lets someone commit to a value in a way that he cannot change it later. The scheme consists of two phases: (i) commit phase, where a commitment is made to a value $v$, but the value is kept secret from third parties; (ii) opening phase, where the committed value $v$ is revealed. A commitment is hiding if an adversary cannot distinguish which of two values corresponds to the commit; it is binding if the committer cannot generate the same commit from two different values. These properties can be either computational, if the adversary is computationally bounded, or perfect, if not. A commitment scheme cannot be both perfectly hiding and perfectly binding, due to fundamental cryptographic limitations (Goldreich, 2004).
Homomorphic commitments, such as Pedersen (Pedersen, 1992) or ElGamal (Elgamal, 1985), preserve the addition or multiplication of corresponding values. In particular, Pedersen commitments are perfectly hiding, computationally binding, and additively homomorphic; ElGamal commitments are computationally hiding, perfectly binding, and multiplicatively homomorphic.
@@ 3.2. Encryption Schemes
Identity-based encryption (IBE):
schemes (Shamir, 1985) are public-key encryption schemes in which the public key is a string that uniquely identifies a user (e.g., an email). The corresponding private keys for decryption are derived from a master private key, by a trusted central authority. IBE can be used to simplify key management and recover, and access control within an organization. Anonymous variants make the public key indistinguishable from the ciphertext, effectively hiding the identity of the recipient of the message.
Threshold encryption:
allows data to be encrypted in such a way that it can only be decrypted when a predefined quorum of participants agrees to reveal it. Threshold encryption can help maintain privacy while removing single point of failures. A popular implementation uses the ElGamal encryption scheme (Elgamal, 1985). A threshold-issuance variant of Boneh-Franklin IBE is used in UTT (Tomescu et al., 2022), to distribute the master secret key among multiple authorities.
Homomorphic encryption:
schemes (Rivest et al., 1978) allow to compute additions, multiplications or both, without revealing the plaintext. Operations can be performed directly on ciphertexts such that the result corresponds to the encryption of the operation applied to the plaintexts. This allows to delegate computation to a third party while maintaining confidentiality. It enables confidential transactions when used as a building block for ZKPs, like to homomorphic commitments (e.g., (Ma et al., 2021; Bünz et al., 2020)).
@@ 3.3. Zero-knowledge Proofs
A ZKP is a protocol enabling a prover to convince a verifier that a statement is true, without disclosing any information beyond the validity of the statement itself. A dishonest prover can convince the verifier of a false statement, only with negligible probability (soundness). If the prover is computationally bounded (computational soundness), the proof is known as argument (Bitansky et al., 2012). ZKPs can be either interactive, requiring the parties to exchange messages, or non-interactive (NIZK). ZKPs can also be proofs of knowledge when the prover can convince the verifier that he knows a certain value, e.g., a secret key (or arguments of knowledge, if the prover is computationally bounded). If the ZK proof is short and fast to verify, it is succinct. If it requires no trusted setup, it is transparent.
A conceptual framework for ZK protocols is the $\Sigma$-protocol, which is a type of 3-move interactive proof system: the prover commits a value, the verifier provides a challenge, and the prover provides a proof based on the commit and the challenge, so that they cannot choose a “malicious” proof. Using the Fiat-Shamir transform, $\Sigma$-protocols can be made non-interactive (Fiat and Shamir, 1987). Building on this, various NIZK protocols have been designed, including zk-SNARK (Bitansky et al., 2012), Bulletproofs (Bünz et al., 2018), and—to some extent—zk-STARK (Ben-Sasson et al., 2019). Zk-SNARKs are succinct non-interactive arguments of knowledge. Many zk-SNARK constructions have been proposed in past years (e.g., (Bitansky et al., 2012; Boneh et al., 2020; Groth, 2016; Chen et al., 2023)), with different approaches also for setup. Among them, the most popular approach is Groth16 (Groth, 2016), an instance optimized for performance but with some trade-offs like requiring a trusted setup. Bulletproof is optimized for range proof and does not require trusted setup. Zk-STARK is a recent approach offering efficient proving times and no need for trusted setup. STARKs are succinct transparent arguments of knowledge. However, in the current state, zk-STARK generates large proof size (a few hundred kilobytes), so they are not as succinct as SNARKs. A recent survey on ZKPs can be found, e.g., in (Christ et al., 2024; El-Hajj and Oude Roelink, 2024).
@@ 3.4. Anonymity-enhanced Signatures
(Chaum, 1983) are a form of digital signature in which the signer produces a valid signature on a message without learning its content: before signing, the message is blinded, i.e., transformed through a random element that makes it unrecognizable. Once the contents of the signed message are unblinded, they can be publicly verified as usual.
Randomizable signatures:
@@
provide a way to randomize a previously generated signature, transforming it into a new valid signature, but making it impossible to link the two instances together. This enables a level of anonymity by unlinking user identities from their transaction history, as each transaction appears distinct even if performed by the same user. The original idea was proposed by Camenisch and Lysyanskaya (Camenisch and Lysyanskaya, 2003, 2004). Later, Pointcheval-Sanders proposed more efficient signature and verification algorithms (Pointcheval and Sanders, 2016). Existing works (e.g., (Baudet et al., 2023; Wüst et al., 2019; Sarencheh et al., 2025; Tomescu et al., 2022)) adopt either Pointcheval-Sanders or Coconut (Sonnino et al., 2020), a threshold scheme derived from it. The Pointcheval-Sanders scheme can produce signatures of a message directly from its commitment; this effectively makes it a blind signature, thanks to the hiding property of commitments.
Group signatures:
@@
schemes (Chaum and van Heyst, 1991) enable any member of a designated group to generate a valid signature without revealing who signed (ensuring anonymity). A verifier uses a group public key to confirm that a valid signature was produced by some group member. This property also implies unlinkability: it is not possible to determine whether two different signatures were produced by the same user (Perera et al., 2022). However, group signatures can be made linkable to prevent double spending in transactions (e.g., (Zhang et al., 2019)). Group signatures also support user traceability: a designated group manager holds a group secret key that enables him to revoke a signer’s anonymity and reveal his identity when necessary.
Ring signatures:
@@
(Rivest et al., 2001) are similar to group signatures, as they hide the signer’s identity behind a group. Differently from group signatures, user anonymity is permanent, because there is no revoke mechanism. Indeed, the signer simply joins their public key with a set of other public keys during the signing phase. A verifier needs the whole list of public keys to verify the signature validity, and cannot discern which public key is of the signer, but can only guess . There are some variant ring signatures, where anonymity is somewhat limited and the signer can be traced under special circumstances, e.g., after double spending. These are called traceable ring signatures  (Fujisaki and Suzuki, 2007; Fujisaki, 2011). Ring signatures can also be linkable (Liu and Wong, 2005), if they allow anyone to determine if two signatures were produced by the same signer (e.g., (Van Saberhagen, 2013; Li et al., 2021)).
Threshold signatures:
@@
are a type of signature that can be jointly produced by a committee of signers. A $t$-of-$n$ signature scheme requires at least $t$ signers out of a total of $n$ participants to produce a valid signature. Works that use threshold signatures (Androulaki et al., 2020; Sarencheh et al., 2025; Kiayias et al., 2022; Baudet et al., 2023) use threshold blind signatures, based on Coconut (Sonnino et al., 2020).

### v3

@@ Abstract.
Crypto-assets and central bank digital currencies (CBDCs) are reshaping how value is exchanged in distributed computing environments. These systems combine cryptographic primitives, protocol design, and system architectures to provide transparency and efficiency while raising critical challenges around privacy and regulatory compliance. This survey offers a comprehensive overview of privacy-preserving digital payment systems, covering both decentralized ledger systems and CBDCs. We present a taxonomy of privacy goals—including anonymity, confidentiality, unlinkability, and auditability—and map them to the cryptographic primitives, protocols, and system architectures that implement them. Our work adopts a design-oriented perspective, linking high-level privacy objectives to concrete implementations. We also trace the evolution of privacy-preserving digital payment systems through three generations, highlighting shifts from basic anonymity guarantees toward more nuanced privacy-accountability trade-offs. Finally, we identify open challenges, motivating further research into architectures and solutions that balance strong privacy with real-world auditability needs.
@@ 1. Introduction
The rapid proliferation of digital payment systems has sparked renewed interest in designing distributed architectures that balance user privacy with institutional accountability. Spanning decentralized crypto-assets [^1] The term cryptocurrency is widely used to describe decentralized systems such as Bitcoin, Ethereum, or Monero, that enable peer-to-peer asset transfer without a central authority. However, this usage differs from the terminology adopted by central banks and regulators (ECB, 2018), which reserve the term “currency” for instruments with legal tender status. To reflect this distinction and bridge the terminology gap between these communities, we adopt the more nuanced terms crypto-assets or digital payment systems throughout the paper. and centrally managed infrastructures such as central bank digital currencies (CBDCs), these systems leverage cryptographic techniques and distributed consensus to enable secure, efficient, and programmable value transfer. However, the same transparency and auditability that make these systems attractive also raise fundamental questions about user privacy, institutional control, and regulatory compliance (e.g., (Auer et al., 2025; Ghesmati et al., 2022; Herskind et al., 2020)). For example, public blockchains typically offer pseudonymity. [^2] Anonymity denotes indistinguishability within a set, while pseudonymity uses identifiers that are not directly linked to real-world identities. This enables transaction tracing, address clustering, and ultimately user de-anonymization, thus exposing financial behavior (e.g., (Meiklejohn et al., 2013; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020; Karame et al., 2015)). As a result, privacy-preserving digital payments have emerged as a critical research field, attracting attention from cryptographers, engineers, economists, and policymakers.
@@
Researchers have proposed a variety of privacy-preserving solutions leveraging privacy-enhancing technologies (PETs), including mixing protocols (e.g., (Maxwell, 2013; Ziegeldorf et al., 2018; Ruffing and Moreno-Sanchez, 2017)), confidential transactions (e.g., (Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022)), and privacy-native crypto-assets (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Tomescu et al., 2022; Noether et al., 2016)). These approaches provide different privacy guarantees, such as anonymity, confidentiality, and unlinkability. However, privacy alone is not sufficient: regulatory frameworks such as Anti-Money Laundering (AML) and Combating the Financing of Terrorism (CFT) require mechanisms for auditability and selective disclosure. To reconcile these conflicting goals, recent work has explored cryptographic tools such as zero-knowledge proofs (ZKPs), threshold encryption, and homomorphic commitments.
@@
This survey proposes a unifying, design-oriented perspective on privacy-preserving digital payment systems, covering both decentralized ledgers and centrally managed systems with privacy guarantees—including CBDCs. It organizes the literature around conceptual dimensions that highlight recurring design choices and trade-offs. Unlike previous surveys that focus on broad overviews (e.g., (Genkin et al., 2018; Auer et al., 2025; Herskind et al., 2020)), specific technologies (e.g., (Almashaqbeh and Solomon, 2022; Ghesmati et al., 2022; Sun et al., 2021)), or do not consider auditability (e.g., (Peng et al., 2021; Herskind et al., 2020; Feng et al., 2019)), our work aims to bridge these dimensions. Our contributions are threefold. First, we introduce a taxonomy that classifies privacy-preserving digital payments systems along multiple dimensions, including system architecture, trust model, value representation, privacy guarantees, and auditability mechanisms. Second, we systematize how privacy properties are formalized and realized, mapping high-level goals to technical mechanisms. We trace the development of privacy-preserving digital payment systems, highlighting the role of cryptographic techniques across three generations. Third, we identify open research challenges stemming from the design of privacy-preserving digital payment systems, which lie at the intersection of cryptography, distributed systems, and privacy engineering. Supplementary material presents the methodology used to collect and select the surveyed literature, and further contextualizes key publication venues and technical details of specific privacy-enhancing mechanisms.
@@ 2. Related Surveys
The rapid growth of crypto-asset systems has spurred multiple efforts to systematize knowledge (SoK) across diverse technical communities. Bonneau et al. (Bonneau et al., 2015) reviewed extensions of Bitcoin and alternative protocols (sometimes referred to as altcoins), including early privacy-native protocols. Genkin et al. (Genkin et al., 2018) offered a broad and accessible overview of techniques for anonymity in decentralized crypto solutions, while emphasizing the absence of a unified definition of privacy. More recent surveys attempted to formalize anonymity notions. For instance, Amarasinghe et al. (Amarasinghe et al., 2019) developed a taxonomy of anonymity features (including anonymity, confidentiality, unlinkability, untraceability, and deniability), and evaluated Bitcoin-derived solutions against it. Our work draws from and extends this analysis by further considering the trade-offs between anonymity, confidentiality, and auditability across a broader range of systems, including CBDCs. Feng et al. (Feng et al., 2019) offered a layered classification of privacy-preserving mechanisms from network-level obfuscation to cryptographic primitives such as ring signatures, Pedersen commitments, homomorphic encryption, and NIZKs. While technically detailed, their coverage of payment protocols and privacy-accountability trade-offs is limited. Similarly, a systematic literature review by Herskind et al. (Herskind et al., 2020) catalogs techniques such as stealth addresses, confidential transactions, and network anonymity. However, it does not analyze core system features such as value representation models or system architectures, which are key components of our taxonomy. Peng et al. (Peng et al., 2021) identify PETs and illustrate how various protocols integrate them, but do not investigate protocol evolution and design trade-offs.
@@
A distinct body of work addresses privacy in CBDCs, driven by the need to balance individual privacy with institutional oversight. Auer et al. (Auer et al., 2022, 2025) distinguish between soft privacy, based on access controls and trust in intermediaries, and hard privacy, based on cryptographic techniques. While they offer a high-level classification and conceptual framework, their discussion does not consider low-level protocol mechanisms. Other regulatory-focused contributions include Darbha and Arora (Bank of Canada) [^3] https://www.bankofcanada.ca/2020/06/staff-analytical-note-2020-9/ and Pocher and Veneris (Pocher and Veneris, 2022) explore PETs from a compliance and governance perspective, mapping privacy mechanisms to AML/CFT auditability requirements in retail CBDC designs. These works offer blueprints but lack detailed cryptographic analysis. Conversely, our work presents a comparative review of technical design choices including primitives, protocols, and architectures.
@@
We propose a unified taxonomy of privacy-preserving payment systems that integrates both decentralized ledger-based solutions and CBDCs. Specifically, we present a systematic analysis of existing solutions to identify key value representation models, privacy features, i.e., anonymity, confidentiality, accountability, and auditability guarantees. We focus on how such features are realized through cryptographic protocols and specific design choices, resulting in a classification of contributions that connects design goals to technical implementations.
@@ 3.1. Commitments and Homomorphic Commitments
A commitment scheme is a cryptographic primitive that allows a party to commit to a value in such a way that it cannot be changed later. The scheme consists of two phases: (i) commit phase, where a commitment is made to a value $v$, but the value is kept secret from third parties; (ii) opening phase, where the committed value $v$ is revealed. A commitment is hiding if an adversary cannot distinguish which of two values corresponds to the commit; it is binding if the committer cannot generate the same commit from two different values. These properties can be either computational, if the adversary is computationally bounded, or perfect, if not. A commitment scheme cannot be both perfectly hiding and perfectly binding, due to fundamental cryptographic limitations (Goldreich, 2004).
Homomorphic commitments preserve the addition or multiplication of corresponding values. For example, let $g$ and $h$ be elements of a prime $q$ group $\mathbb{G}_{q}$ such that $\log_{g}h$ is not feasible to compute, let $r_{i}$ be a random value over $\mathbb{Z}_{q}$ and let $v_{i}\in\mathbb{Z}_{q}$ be the committer’s value. Perdersen commitments are defined as $c_{v_{i}}=g^{v_{i}}h^{r_{i}}$. The homomorphic properties of Pedersen commitments guarantee that $g^{v_{1}}h^{r_{1}}\cdot g^{v_{2}}h^{r_{2}}=g^{v_{1}+v_{2}}h^{r_{1}+r_{2}}$ namely, the sum of the commitments of $v_{1}$ and $v_{2}$ is equal to the commitment of the sum of the two values under the sum of the two randomness $r_{1}$ and $r_{2}$, allowing multiple operations on the committed messages. In particular, Pedersen (Pedersen, 1992) commitments are perfectly hiding, computationally binding, and additively homomorphic; ElGamal (Elgamal, 1985) commitments are computationally hiding, perfectly binding, and multiplicatively homomorphic.
@@ 3.2. Encryption Schemes
Identity-based encryption (IBE)
This scheme (Shamir, 1985) is a public-key encryption scheme in which the public key is a string that uniquely identifies a user (e.g., an email). The corresponding private keys for decryption are derived from a master private key, by a trusted central authority. IBE can be used to simplify key management and recover, and access control within an organization. Anonymous variants make the public key indistinguishable from the ciphertext, effectively hiding the identity of the recipient of the message.
Threshold encryption
It allows data to be encrypted in such a way that it can only be decrypted when a predefined quorum of participants agrees to reveal it. Threshold encryption can help maintain privacy while removing single point of failures. A popular implementation uses the ElGamal encryption scheme (Elgamal, 1985). A threshold-issuance variant of Boneh-Franklin IBE is used in UTT (Tomescu et al., 2022), to distribute the master secret key among multiple authorities.
Homomorphic encryption
This technique (Rivest et al., 1978) allows computing additions, multiplications or both, without revealing the plaintext. Operations can be performed directly on ciphertexts such that the result corresponds to the encryption of the same operation applied to the plaintexts, i.e., $\mathrm{Enc}(m_{1})\star\mathrm{Enc}(m_{2})=\mathrm{Enc}(m_{1}\star m_{2})$, where $\mathrm{Enc}(\cdot)$ represent an encryption function and $\star$ denotes either addition or multiplication. This allows to delegate computation to a third party while maintaining confidentiality. It enables confidential transactions when used as a building block for ZKPs, like to homomorphic commitments (e.g., (Ma et al., 2021; Bünz et al., 2020)).
@@ 3.3. Zero-knowledge Proofs
A ZKP is a protocol enabling a prover to convince a verifier that a statement is true, without disclosing any information beyond the validity of the statement itself. A dishonest prover can convince the verifier of a false statement, only with negligible probability (soundness). If the prover is computationally bounded (computational soundness), the proof is known as argument (Bitansky et al., 2012). ZKPs can be either interactive, requiring the parties to exchange messages, or non-interactive (NIZK), when the verifier is convinced through a single message. ZKPs can also be proofs of knowledge when the prover can convince the verifier that he knows a certain value, e.g., a secret key (or arguments of knowledge, if the prover is computationally bounded). If the ZKP is short and fast to verify, it is said to be succinct. If it requires no trusted setup, it is transparent.
A conceptual framework for ZK protocols is the $\Sigma$-protocol, which is a type of 3-move interactive proof system: the prover commits a value, the verifier provides a challenge, and the prover provides a proof based on the commit and the challenge, so that they cannot choose a “malicious” proof. Using the Fiat-Shamir transform, $\Sigma$-protocols can be made non-interactive (Fiat and Shamir, 1987). Building on this, various NIZK protocols have been designed, including zk-SNARK (Bitansky et al., 2012), Bulletproofs (Bünz et al., 2018), and—to some extent—zk-STARK (Ben-Sasson et al., 2019). Zk-SNARKs are succinct non-interactive arguments of knowledge. Many zk-SNARK constructions have been proposed in past years (e.g., (Bitansky et al., 2012; Boneh et al., 2020; Groth, 2016; Chen et al., 2023)), with different approaches also for setup. Among them, the most popular approach is Groth16 (Groth, 2016), an instance optimized for performance but with some trade-offs like requiring a trusted setup. Bulletproofs are optimized for range proof and does not require trusted setup. Zk-STARK is a recent approach offering efficient proving times and no need for trusted setup. STARKs are succinct transparent arguments of knowledge. However, in the current state, zk-STARK generates large proof size (a few hundred kilobytes), so they are not as succinct as SNARKs. A recent survey on ZKPs can be found, e.g., in (Christ et al., 2024; El-Hajj and Oude Roelink, 2024).
@@ 3.4. Anonymity-enhanced Signatures
Blind signatures (Chaum, 1983) are a form of digital signature in which the signer produces a valid signature on a message without learning its content: before signing, the message is blinded, i.e., transformed through a random element that makes it unrecognizable. Once the contents of the signed message are unblinded, they can be publicly verified as usual.
Randomizable signatures
@@
They provide a way to randomize a previously generated signature, transforming it into a new valid signature, but making it impossible to link the two instances together. This enables a level of anonymity by unlinking user identities from their transaction history, as each transaction appears distinct even if performed by the same user. The original idea was proposed by Camenisch and Lysyanskaya (Camenisch and Lysyanskaya, 2003, 2004) but suffered from a linear size in the number of messages to be signed. Later, Pointcheval-Sanders proposed more efficient signature and verification algorithms (Pointcheval and Sanders, 2016). Existing works (e.g., (Baudet et al., 2023; Wüst et al., 2019; Sarencheh et al., 2025; Tomescu et al., 2022)) adopt either Pointcheval-Sanders or Coconut (Sonnino et al., 2020), a threshold scheme derived from it. The Pointcheval-Sanders scheme can produce signatures of a message directly from its commitment; this effectively makes it a blind signature, thanks to the hiding property of commitments.
Group signatures
@@
This scheme (Chaum and van Heyst, 1991) enables any member of a designated group to generate a valid signature without revealing who signed (ensuring anonymity). A verifier uses a group public key to confirm that a valid signature was produced by some group member. This property also implies unlinkability: it is not possible to determine whether two different signatures were produced by the same user (Perera et al., 2022). However, group signatures can be made linkable to prevent double spending in transactions (e.g., (Zhang et al., 2019)). Group signatures also support user traceability: a designated group manager holds a group secret key that enables him to revoke a signer’s anonymity and reveal his identity when necessary.
Ring signatures
@@
Ring signatures (Rivest et al., 2001) are similar to group signatures, as they hide the signer’s identity behind a group. Differently from group signatures, user anonymity is permanent, because there is no revoke mechanism. Indeed, the signer simply joins their public key with a set of other public keys during the signing phase. A verifier needs the whole list of public keys to verify the signature validity, and cannot discern which of the $n$ public keys is of the signer, but can only guess with probability $1/n$. There are some variant ring signatures, where anonymity is somewhat limited and the signer can be traced under special circumstances, e.g., after double spending. These are called traceable ring signatures  (Fujisaki and Suzuki, 2007; Fujisaki, 2011). Ring signatures can also be linkable (Liu and Wong, 2005), if they allow anyone to determine if two signatures were produced by the same signer (e.g., (Van Saberhagen, 2013; Li et al., 2021)).
Threshold signatures
@@
They are a type of signature that can be jointly produced by a committee of signers. A $t$-of-$n$ signature scheme requires at least $t$ signers out of a total of $n$ participants to produce a valid signature. Works that use threshold signatures (Androulaki et al., 2020; Sarencheh et al., 2025; Kiayias et al., 2022; Baudet et al., 2023) use threshold blind signatures, based on Coconut (Sonnino et al., 2020).